# Contrary Inferences for Classical Histories within the Consistent Histories Formulation of Quantum Theory

### v3

@@ 1 Introduction
The consistent histories formulation of quantum theory [4, 5, 6, 7, 8, 1] brings forward a proposal for dealing with the quantum mechanics of a closed system by replacing the observer-dependent process of quantum measurement with a consistency condition. It is this condition that determines when a set of (coarse-grained) histories admits a classical description. However, the formalism allows for multiple consistent sets to exist for the same physical situation and the dominant view is that all these sets are to be treated equally [^1] An alternative view is that one consistent set (or a family of consistent sets) are more natural (in some suitably defined sense) and should be given a preferred status in interpreting the theory. The physical principle that selects such consistent set(s) is still elusive even though certain proposals have been given [9, 10, 11]. Adrian Kent in ref. [3] noted that, if we were to combine propositions belonging to different consistent sets, we would reach to “paradoxes” in the emergent classical logic such as contrary inferences. This reinforced the view, shared by proponents of the approach, that each consistent set is a context and that one should reason within a single consistent set each time. This would avoid the aforementioned paradoxes [12]. Contextuality, that is, reaching to different conclusions depending on the context in which the question is asked as well as the subsequent “paradoxes” that appear when one mixes contexts is a well-known and fundamental property of the (microscopic) quantum systems [13, 14, 15, 16]. It is also true though that such paradoxes should not persist in larger scales when dealing, for example, with quasi-classical systems that have already decohered in some quasi-classical basis (e.g. the overcomplete basis of localized wavepackets) when subjected to local measurements.
@@
where $V_{\bar{\Delta}}$ is the potential of the infinite square well, see eq. (6). Now, in order for this partition to be a consistent set, the two cells of the histories space need to have zero interference. As was shown in [17], this leads to the condition $\bra{\Psi(\tau)}\Psi_{r}(\tau)\rangle=1$ together with a boundary condition, where |$\ket{\Psi(\tau)}$ is the system unitarily evolved until the final time $\tau$ and |$\ket{\Psi_{r}(\tau)}$ is the evolution of the system with the restriction that it did not leave the cell we ask for, see eq. (2) for a derivation. It is very important to note that in the consistent histories formulation there is no actual physical measurement taking place – we are dealing with closed quantum systems without any observer – and the “barriers” that are mathematically represented by projection operators is just a simpler way to describe the corresponding path integral that will tell us if the consistency condition is satisfied (and thus we can assign classical probabilities) or not (and thus this question involves quantumness and we cannot assign classical probabilities to those histories). Now, it is known that there are more than one consistent sets for the same histories space which means that we can define classical probabilities in different consistent sets and potentially get different predictions (probabilities) in each case. This happens when we have (at least) two different partitions of the histories space and when some coarse-grained histories from each partition form a zero measure cover of the histories space. Then it was shown in [10] that these propositions are contrary. In the case they both have probability one to happen this constitutes a problem for the quantum theory which becomes severe when we deal with the quasi-classical limit as we discuss in this work.
@@
    where $C_{\Delta}$ is the set of histories that stay in $\Delta$.
@@
    where $\rho$ is the initial density matrix.
@@ 2 The time-of-arrival problem in the consistent histories approach and contrary inferences
The question “what is the probability that the particle will be (“arrive”) in some region (range) $\Delta$ during some interval $[0,\tau]$?” is usually referred to as the time-of-arrival problem. The set-up of the problem, shown in Figure 1(a), is that of a free particle with support in the interval $\Delta$ and its complement $\bar{\Delta}$ defined respectively by $\Delta=[0,+\infty)$ and $\bar{\Delta}=(-\infty,0]$, which evolves under the action of the free one-dimensional Hamiltonian. The particle is initially located in the negative axis with positive average momentum and it is thus expected to cross the surface $x=0$ at some moment $t\in[0,\tau]$. We consider histories that frequently check if the particle is still in $\bar{\Delta}$, depicted in red colour in Figure 1(b), by applying very frequently the continuous projection operator
[…]
@@
and defined on the proper subspace $L^{2}(\bar{\Delta})$ with domain $D(H_{r})=\{\Psi\in\mathcal{L}^{2}(\mathbb{R})|\partial_{x}\Psi\in\mathcal{L}^{2}(\mathbb{R}),\Psi(\partial\bar{\Delta}=0)\}$. This can alternatively be written as [39]
@@
which means that the particle behaves as if it were confined in $\bar{\Delta}$ by rigid barriers when one makes infinitely many measurements (projections with (3)). [^3] We will call the walls induced by infinitely many projections “barriers” while we keep the term “wall” for the boundaries of the infinite square well in which all the dynamics happens. This construction has also been called “Zeno box” [44], because it is the analogue of the quantum Zeno effect for continuous variables systems. We emphasise that this is only a mathematical equivalence, since in our considerations there are no measurements, the projections arise in our effort to compute the amplitude of a particular (restricted) set of histories.
@@
$D(h,\bar{h})=0$ $=$ $\bra{\Psi_{0}}g^{\dagger}_{r}(t,t_{0})\left(g(t,t_{0})-g_{r}(t,t_{0})\right)\ket{\Psi_{0}}$
$=$ $\bra{\Psi_{0}}g^{\dagger}_{r}(t,t_{0})g(t,t_{0})\ket{\Psi_{0}}-\bra{\Psi_{0}}g^{\dagger}_{r}(t,t_{0})g_{r}(t,t_{0})\ket{\Psi_{0}}$
$=$ $\bra{\Psi_{0}}g^{\dagger}_{r}(t,t_{0})g(t,t_{0})\ket{\Psi_{0}}-\bra{\Psi_{0}}\bar{P}\ket{\Psi_{0}}$
$\bra{\Psi_{0}}\bar{P}\ket{\Psi_{0}}$ $=$ $\bra{\Psi_{0}}g^{\dagger}_{r}(t,t_{0})g(t,t_{0})\ket{\Psi_{0}}.$ (9)
where we have used that i) $g_{r}(t,t_{0})\ket{\Psi_{0}}$ is the evolution restricted in $\bar{\Delta}$ and corresponds to history $h$, ii) $\left(g(t,t_{0})-g_{r}(t,t_{0})\right)\ket{\Psi_{0}}$ is the evolution that is not always in the region $\bar{\Delta}$ and corresponds to the history $\bar{h}$ and iii) if |$\ket{\Psi_{0}}$ is initialised in the region $\bar{\Delta}$ then we have $\bra{\Psi_{0}}\bar{P}\ket{\Psi_{0}}=1$. This relation, with the above assumptions therefore gives:
$\bra{\Psi_{0}}g_{r}^{\dagger}(t,t_{0})g(t,t_{0})\ket{\Psi_{0}}=\bra{\Psi_{r}(t)}\ket{\Psi(t)}=1$ (10)
@@
$\innerproduct{x}{\Psi_{0}}_{\partial\bar{\Delta}}=0$ (12)
These are the conditions we will later use to construct our consistent sets and to show that they lead to contrary inferences. This is also possible because we can identify coarse-grained histories with propositions (and operators). Two propositions represented by the projectors $P,\ Q$ at some moment of time are called contrary if $PQ=QP=0$ and $P<1-Q$, [^5] We can write $P_{A}P_{B}=P_{A}\wedge P_{B}:=P_{A\wedge B}$ only when $[P_{A},P_{B}]=0$. If in addition $P_{A}P_{B}=P_{B}P_{A}=0$ we can write for the conjunction $P_{A\vee B}:=P_{A}\vee P_{B}=P_{A}+P_{B}$, see e.g. [45]. i.e. if they do not intersect and jointly do not span the full space at that moment. [^6] Compare with the definition of the contradictory propositions, which are those who do not intersect but cover the full space jointly at that moment, i.e. $PQ=QP=0$ and $P=1-Q$.
A proposition that happens with probability one is assigned truth value True. It is not hard to see that if one has two contrary propositions (say $P,Q$) that both occur with probability one, it means that they both have truth value $v(P)=v(Q)=\textrm{True}$, something that leads to logical paradoxes. By the law of inference (modus ponens), if a proposition $P$ is assigned truth value True, so does any proposition that contains $P$. However, we also have that $1-Q>P$ so that the proposition $1-Q$ should also take value True which leads to the situation that both $v(Q)$ and the negation $v(1-Q)$ are True. To reach this paradox we had to make an inference (thus the term “contrary inferences”). Interestingly, in consistent histories, one can get two contrary propositions having probability one (each defined in a different consistent set).
@@ 3.1 Quasi-classical and macroscopic approximation
Figure 2: Probability distribution of the coherent state in position and momentum space at $t=0$. The parameters were chosen as follows: $n_{0}=201$ and $\sigma_{0}=20$. The phase $\varphi_{0}$ is selected such that the wave function is localised at $x/L=-1/2$, as demonstrated in the figure.
In this section we will show that there exist (at least) two different consistent sets, each of which gives a different answer to the time-of-arrival question. We explicitly construct them and show that they satisfy the consistency conditions (10) and (12). We consider a coherent state to represent the particle in an infinite square well $[-L,L]$ with Hamiltonian
$H=-\frac{\hbar^{2}}{2M}\frac{d^{2}}{dx^{2}}+V(x),\quad V(x)=\begin{cases}0,&\absolutevalue{x}<L,\\ \infty,&\absolutevalue{x}\geq L\end{cases}$ (16)
where $M$ denotes the mass of the particle and $\hbar$ the reduced Planck constant. This Hamiltonian is a self-adjoint operator on the vector space $\mathcal{L}^{2}(\Delta_{2L})$ and describes the free evolution in $\Delta_{2L}=[-L,L]$. To construct the coherent state, we start with a wave function of the form
@@
Since we have justified that our system describes a quasi-classical and macroscopic system, we would not expect to observe any of the quantum paradoxes that arise when considering microscopic systems. We proceed by giving the two different coarse-grainings that, within the consistent histories formulation, lead to contrary inferences for this quasi-classical and macroscopic system. Specifically, we pose the question of the time-of-arrival problem for a quasi-classical particle represented as a suitably localised wave packet initially located at $x=-\frac{L}{2}$ with positive average momentum $p=p_{0}$ and confined in an infinite square well $[-L,L]$. We deal with the question on whether it passes $x=0$ or not by considering the complementary question of whether the particle always remains in the subregion $[-L,0]$ and we construct “Zeno” histories [36].
@@ 3.2 Construction of the coarse-grainings: The first consistent set
Figure 3: Left Panel: The time-of-arrival problem in the infinite square well; this is an instant of the partition for the first set of histories. The two subregions are $\Delta_{1}=[0,L]$ and $\bar{\Delta}_{1}=[-L,0]$. The question is whether the particle, equipped with positive momentum will be even for only one moment in $\Delta_{1}$. The barrier at $x=0$ is fictitious and corresponds to infinitely many projections with the operator (38). The projection takes place also at $x=-L$ but this also coincides with the real barrier of the square well.
@@
It is important to note here, that our conclusion is valid even if the consistency condition is satisfied for a single specific choice of $\tau$. In the following we will treat $\tau$ as variable and see for which choices of $\tau$ the condition is satisfied. Interestingly, we will find that the consistency condition is satisfied for large “intervals” so it is “robust” in time. For example a good choice for our aims is $\tau=0.85T_{cl}$, in units of the classical period of the well $[-L,L]$, when the centre of the wave packet has returned to its initial position but with opposite average momentum.
@@
To show that this set is consistent, we need to check the relation (10) and the boundary condition (12). For the boundary condition to be satisfied we must have $\Psi(-L,0)=0$. This holds trivially since it is one of the boundaries of the $[-L,L]$, and we have caclculated that $\Psi(0,0)=0$ up to the order $O(10^{-25})$. For the condition (10) to hold we have to find at which times the overlap
@@
Figure 4: The square of the overlap of the two wave packets as a function of time normalised with the classical time (right plot) and the "zoomed" plot near $\tau=0.85T_{cl}$ (right plot). The parameters the same as follows $n_{0}^{1}=100$, $n_{0}=2n_{0}^{1}+1$, $\sigma_{0}^{1}=10$, $\sigma_{0}=2\sigma_{0}^{1}$. The phases $\varphi_{0}^{1}$ and $\varphi_{0}$ are selected such that both wave functions are localised at $x/L=-1/2$. The superscript $1$ refers to the coherent state in the well spanning $[-L,0]$.
Probabilities
The candidate probability, for remaining during all the interval $[0,\tau]$ in the region $\bar{\Delta}_{1}$ is given by the expression $D(h_{1},h_{1})=\mu(h_{1})$ and it is not difficult to see that it is one for $\tau=0.85T_{cl}$:
$\mu(h_{1})=\bra{\Psi_{0}}g_{r}^{\dagger}(\tau,0)g_{r}(\tau,0)\ket{\Psi_{0}}\approx 1$ (25)
@@
$\mu(\bar{h}_{1})$ $=$ $\bra{\Psi_{0}}(g^{\dagger}(\tau,0)-g_{r}^{\dagger}(\tau,0))\times(g(\tau,0)-g_{r}(\tau,0)\ket{\Psi_{0}}$ (26)
$=$ $\bra{\Psi_{0}}g^{\dagger}(\tau,0)g(\tau,0)\ket{\Psi_{0}}-\bra{\Psi_{0}}g_{r}^{\dagger}(\tau,0)g(\tau,0)\ket{\Psi_{0}}-$
$-\bra{\Psi_{0}}g^{\dagger}(\tau,0)g_{r}(\tau,0)\ket{\Psi_{0}}+\bra{\Psi_{0}}g_{r}^{\dagger}(\tau,0)g_{r}(\tau,0)\ket{\Psi_{0}}$
@@
These show that the answer to the initial question of the particle crossing the origin within the time interval $[0,\tau]$, is negative.
@@ 3.3 Construction of the coarse-grainings: The second consistent set (time-dependent)
Here we are interested in histories that remain within a (time-dependent) interval that at all times contains the centre of the wave packet and is sufficiently wide to include essentially the full wave packet. One can define such a time-dependent interval in different ways and care needs to be taken when the classical trajectory (and the centre of the wave packet) hits one of the walls placed at ±$\pm L$. The first important point is that this interval should be much larger than the width of the wave packet at all times, and this can be achieved if the moving interval is $O(L)$ since the length of the well $L$ is significantly larger than the width of the Gaussian wavepacket in the position space which in our case it is true. The approach we will take is that at each moment of time we will apply the position projection operator
[…]
@@
We take the interval $\Delta_{2}(t)$ to have width $\lambda=L/8$ and the mathematical effect of these projections is to introduce “fictitious” infinite potential walls at the boundary of this region. It is clear that when $x_{cl}(t)+L/8>L$ or $x_{cl}-L/8<-L$, the fictitious wall is further than the real walls that exist in ±$\pm L$. This means that the box (infinite square well) that the restricted evolution takes place, in those cases, changes size to reach $L/8$ when $x_{cl}(t)=L$ from the initial $L/4$ width. Since the length of the well is significantly larger than the width of the Gaussian wavepacket this does not change significantly the analysis. Alternatively, one could use a different interval that “freezes” when $x_{cl}(t_{1})+L/8=L$ and then starts following the wave packet at $t_{2}$ when the centre has returned at $87L/$ 8 but is moving the opposite direction. We focus on the first case, but the consistency condition and probabilities would be the same in the second case too.
The interval $[-L,L]$ is divided in three subregions; we are interested in the histories in the second one, those that never leave $\Delta_{2}$. These are constructed by all the paths reflected from the boundaries of the interval. In this way, the set of histories belonging in the interval $\Delta_{2}(t)$ are:
@@
It can be seen that at the initial moment $t_{0}=0$, the overlap between the wave packet defined in $\Delta_{2L}$ and its restriction in $[x_{cl}-\lambda,x_{cl}+\lambda]$ is equal to one:
[…]
because we have selected the interval $\Delta_{2}(t_{0})$. The translational symmetry guarantees that this value will be approximately preserved in time while it moves in the infinite square well because of the quasi-classical properties of the Gaussian state. Viewing this differently, the unrestricted evolution of our quasi-classical system does not spread and keeps the same width in position and momentum space, while the centre of the wave packet follows the classical trajectory. By construction the same holds for the restricted evolution. Moreover, the wave packet bounces at the walls ±$\pm L$ in both cases, while the fictitious walls of the restricted evolution are always at point that the wave function is vanishingly small because $L/8\gg$ is much larger than the width of the wave packet.
@@
What is crucial to note here is that all the paths that remained through all the interval $[0,\tau]$ in $\Delta_{2}(t)$ passed through $x=0$. This is true since, for example, at $t=T_{cl}/4$, the interval $\Delta_{2}(t)=[3L/8,5L/8]$ and is fully in the positive axis meaning that paths that at that moment are within $\Delta_{2}(t)$ have been in the $x>0$ region. In other words $h_{2}$ is a strict subset of the histories that crossed $x=0$. We can therefore infer from this consistent set, that the particle passed $x=0$ with certainty (probability unity). Interestingly, while this consistent set gives an indirect way to answer the time-of-arrival question, it actually recovers the answer that one would intuitively expect.

### v4

@@ 1 Introduction
The consistent histories formulation of quantum theory [4, 5, 6, 7, 8, 1] brings forward a proposal for dealing with the quantum mechanics of a closed system by replacing the observer-dependent process of quantum measurement with a consistency condition. It is this condition that determines when a set of (coarse-grained) histories admits a classical description. However, the formalism allows for multiple consistent sets to exist for the same physical situation and the dominant view is that all these sets are to be treated equally [^1] An alternative view is that one consistent set (or a family of consistent sets) are more natural (in some suitably defined sense) and should be given a preferred status in interpreting the theory. The physical principle that selects such consistent set(s) is still elusive even though certain proposals have been given [9, 10, 11]. Adrian Kent in ref. [3] noted that, if we were to combine propositions belonging to different consistent sets, we would reach to “paradoxes” in the emergent classical logic such as contrary inferences. Two propositions represented by the projectors P,QP,\ Q at some moment of time are called contrary if $PQ=QP=0$ and P<1QP<1-Q, [^2] We can write $P_{A}P_{B}=P_{A}\wedge P_{B}:=P_{A\wedge B}$ only when $P_{A},P_{B}]=0$. If in addition $P_{A}P_{B}=P_{B}P_{A}=0$ we can write for the conjunction $P_{A\vee B}:=P_{A}\vee P_{B}=P_{A}+P_{B}$, see e.g. [12]. i.e. if they do not intersect and jointly do not span the full space at that moment. [^3] Compare with the definition of the contradictory propositions, which are those who do not intersect but cover the full space jointly at that moment, i.e. $PQ=QP=0$ and P=1QP=1-Q. This reinforced the view, shared by proponents of the approach, that each consistent set is a context and that one should reason within a single consistent set each time. This would avoid the aforementioned paradoxes [13]. Contextuality, that is, reaching to different conclusions depending on the context in which the question is asked as well as the subsequent “paradoxes” that appear when one mixes contexts is a well-known and fundamental property of the (microscopic) quantum systems [14, 15, 16, 17]. It is also true though that such paradoxes should not persist in larger scales when dealing, for example, with quasi-classical systems that have already decohered in some quasi-classical basis (e.g. the overcomplete basis of localized wavepackets) when subjected to local measurements.
@@
[…]
@@
    where $C_{\Delta}$ is the class operator corresponding to the set of histories that stay in $\Delta$.
@@
    where $\rho$ is the initial density matrix and $\psi_{r}$ the state evolving with the restricted propagator.
@@ 2 The time-of-arrival problem in the consistent histories approach and contrary inferences
The question “what is the probability that the particle will be (“arrive”) in some region (range) $\Delta$ during some interval $0,\tau]$?” is usually referred to as the time-of-arrival problem. The set-up of the problem, shown in Figure 1(a), is that of a free particle with support in the interval $\Delta$ and its complement $\bar{\Delta}$ defined respectively by $\Delta=[0,+\infty)$ and $\bar{\Delta}=(-\infty,0]$, which evolves under the action of the free one-dimensional Hamiltonian. The particle is initially located in the negative axis with positive average momentum and it is thus expected to cross the surface $x=0$ at some moment $t\in[0,\tau]$. We consider histories that frequently check if the particle is still in $\bar{\Delta}$, depicted in red colour in Figure 1(b), by applying very frequently the continuous projection operator
$|.\bar{P}=\int_{\bar{\Delta}}dx\outerproduct{x}{x}$. (3)
@@
and defined on the proper subspace $\mathcal{L}^{2}(\bar{\Delta})$ with domain $D(H_{r})=\{\Psi\in\mathcal{L}^{2}(\mathbb{R})|\partial_{x}\Psi\in\mathcal{L}^{2}(\mathbb{R}),\Psi(\partial\bar{\Delta}=0)\}$. This can alternatively be written as [40]
@@
where $M$ denotes the mass of the particle and $\hbar$ the reduced Planck constant, which means that the particle behaves as if it were confined in $\bar{\Delta}$ by rigid barriers when one makes infinitely many measurements (projections with (3)). [^5] We will call the walls induced by infinitely many projections “barriers” while we keep the term “wall” for the boundaries of the infinite square well in which all the dynamics happens. This construction has also been called “Zeno box” [45], because it is the analogue of the quantum Zeno effect for continuous variables systems. We emphasise that this is only a mathematical equivalence, since in our considerations there are no measurements, the projections arise in our effort to compute the amplitude of a particular (restricted) set of histories.
@@
[…]
@@
[…]
These are the conditions we will later use to construct our consistent sets and to show that they lead to contrary inferences. This is also possible because we can identify coarse-grained histories with propositions (and operators).
A proposition that happens with probability one is assigned truth value True. It is not hard to see that if one has two contrary propositions (say $P,Q$) that both occur with probability one, it means that they both have truth value $v(P)=v(Q)=\textrm{True}$, something that leads to logical paradoxes. By the law of inference (modus ponens), if a proposition $P$ is assigned truth value True, so does any proposition that contains $P$. However, we also have that 1Q>P1-Q>P so that the proposition 1Q1-Q should also take value True which leads to the situation that both $v(Q)$ and the negation v(1Q)v(1-Q) are True. To reach this paradox we had to make an inference (thus the term “contrary inferences”). Interestingly, in consistent histories, one can get two contrary propositions having probability one (each defined in a different consistent set).
@@ 3.1 Quasi-classical and macroscopic approximation
Figure 2: Probability distribution of the coherent state in position and momentum space at $t=0$. The parameters were chosen as follows: $n_{0}=201$ and $\sigma_{0}=20$. The phase $\varphi_{0}$ is selected such that the wave function is localised at x/L=1/2x/L=-1/2, as demonstrated in the figure.
In this section we will show that there exist (at least) two different consistent sets, each of which gives a different answer to the time-of-arrival question. We explicitly construct them and show that they satisfy the consistency conditions (10) and (12). We consider a coherent state to represent the particle in an infinite square well [L,L][-L,L] with Hamiltonian
[…]
This Hamiltonian is a self-adjoint operator on the vector space $\mathcal{L}^{2}(\Delta_{2L})$ and describes the free evolution in $\Delta_{2L}=[-L,L]$. To construct the coherent state, we start with a wave function of the form
@@
Since we have justified that our system describes a quasi-classical and macroscopic system, we would not expect to observe any of the quantum paradoxes that arise when considering microscopic systems. We proceed by giving the two different coarse-grainings that, within the consistent histories formulation, lead to contrary inferences for this quasi-classical and macroscopic system. Specifically, we pose the question of the time-of-arrival problem for a quasi-classical particle represented as a suitably localised wave packet initially located at $x=-\frac{L}{2}$ with positive average momentum $p=p_{0}$ and confined in an infinite square well [L,L][-L,L]. We deal with the question on whether it passes $x=0$ or not by considering the complementary question of whether the particle always remains in the subregion [L,0][-L,0] and we construct “Zeno” histories [37].
@@ 3.2 Construction of the coarse-grainings: The first consistent set
Figure 3: Left Panel: The time-of-arrival problem in the infinite square well; this is an instant of the partition for the first set of histories. The two subregions are $\Delta_{1}=[0,L]$ and $\bar{\Delta}_{1}=[-L,0]$. The question is whether the particle, equipped with positive momentum will be even for only one moment in $\Delta_{1}$. The barrier at $x=0$ is fictitious and corresponds to infinitely many projections with the operator (38). The projection takes place also at x=Lx=-L but this also coincides with the real barrier of the square well.
@@
It is important to note here, that our conclusion is valid even if the consistency condition is satisfied for a single specific choice of $\tau$. In the following we will treat $\tau$ as variable and see for which choices of $\tau$ the condition is satisfied. Interestingly, we will find that the consistency condition is satisfied for large “intervals” so it is “robust” in time. For example a good choice for our aims is $\tau=0.85T_{cl}$, in units of the classical period of the well [L,L][-L,L], when the centre of the wave packet has returned to its initial position but with opposite average momentum.
@@
To show that this set is consistent, we need to check the relation (10) and the boundary condition (12). For the boundary condition to be satisfied we must have $\Psi(-L,0)=0$. This holds trivially since it is one of the boundaries of the [L,L][-L,L], and we have caclculated that $\Psi(0,0)=0$ up to the order $O(10^{-25})$. For the condition (10) to hold we have to find at which times the overlap
$A(t)=\langle\bar{\Psi}_{1}(x,t)\Psi(x,t)\rangle=\int_{-L}^{0}{dx}\bar{\Psi}_{1}(x,t)\Psi(x,t)$ (24)
Figure 4: The square of the overlap of the two wave packets as a function of time normalised with the classical time (left plot) and the "zoomed" plot near $\tau=0.85T_{cl}$ (right plot). The parameters the same as follows $n_{0}^{1}=100$, $n_{0}=2n_{0}^{1}+1$, $\sigma_{0}^{1}=10$, $\sigma_{0}=2\sigma_{0}^{1}$. The phases $\varphi_{0}^{1}$ and $\varphi_{0}$ are selected such that both wave functions are localised at x/L=1/2x/L=-1/2. The superscript 11 refers to the coherent state in the well spanning [L,0][-L,0].
Probabilities
The candidate probability, for remaining during all the interval $0,\tau]$ in the region $\bar{\Delta}_{1}$ is given by the expression $D(h_{1},h_{1})=\mu(h_{1})$ and it is not difficult to see that it is one for $\tau=0.85T_{cl}$:
@@
[…]
@@
These show that the answer to the initial question of the particle crossing the origin within the time interval $0,\tau]$, is negative.
@@ 3.3 Construction of the coarse-grainings: The second consistent set (time-dependent)
Here we are interested in histories that remain within a (time-dependent) interval that at all times contains the centre of the wave packet and is sufficiently wide to include essentially the full wave packet. One can define such a time-dependent interval in different ways and care needs to be taken when the classical trajectory (and the centre of the wave packet) hits one of the walls placed at $\pm L$. The first important point is that this interval should be much larger than the width of the wave packet at all times, and this can be achieved if the moving interval is $O(L)$ since the length of the well $L$ is significantly larger than the width of the Gaussian wavepacket in the position space which in our case it is true. The approach we will take is that at each moment of time we will apply the position projection operator
$|P_{2}(t)=\int_{\Delta_{2}(t)}dx\outerproduct{x}{x}$ (28)
@@
We take the interval $\Delta_{2}(t)$ to have width $\lambda=L/8$ and the mathematical effect of these projections is to introduce “fictitious” infinite potential walls at the boundary of this region. It is clear that when $x_{cl}(t)+L/8>L$ or $x_{cl}-L/8<-L$, the fictitious wall is further than the real walls that exist in $\pm L$. This means that the box (infinite square well) that the restricted evolution takes place, in those cases, changes size to reach $L/8$ when $x_{cl}(t)=L$ from the initial $L/4$ width. Since the length of the well is significantly larger than the width of the Gaussian wavepacket this does not change significantly the analysis. Alternatively, one could use a different interval that “freezes” when $x_{cl}(t_{1})+L/8=L$ and then starts following the wave packet at $t_{2}$ when the centre has returned at $7L/8$ but is moving the opposite direction. We focus on the first case, but the consistency condition and probabilities would be the same in the second case too.
The interval [L,L][-L,L] is divided in three subregions; we are interested in the histories in the second one, those that never leave $\Delta_{2}$. These are constructed by all the paths reflected from the boundaries of the interval. In this way, the set of histories belonging in the interval $\Delta_{2}(t)$ are:
@@
It can be seen that at the initial moment $t_{0}=0$, the overlap between the wave packet defined in $\Delta_{2L}$ and its restriction in $x_{cl}-\lambda,x_{cl}+\lambda]$ is equal to one:
$\int_{x_{cl}-\lambda}^{x_{cl}+\lambda}\Psi^{*}(x,t_{0})\Psi_{2}(x,t_{0})dx\approx 1$ (30)
because we have selected the interval $\Delta_{2}(t_{0})$. The translational symmetry guarantees that this value will be approximately preserved in time while it moves in the infinite square well because of the quasi-classical properties of the Gaussian state. Viewing this differently, the unrestricted evolution of our quasi-classical system does not spread and keeps the same width in position and momentum space, while the centre of the wave packet follows the classical trajectory. By construction the same holds for the restricted evolution. Moreover, the wave packet bounces at the walls $\pm L$ in both cases, while the fictitious walls of the restricted evolution are always at point that the wave function is vanishingly small because $L/8\gg$ is much larger than the width of the wave packet.
@@
What is crucial to note here is that all the paths that remained through all the interval $0,\tau]$ in $\Delta_{2}(t)$ passed through $x=0$. This is true since, for example, at $t=T_{cl}/4$, the interval $\Delta_{2}(t)=[3L/8,5L/8]$ and is fully in the positive axis meaning that paths that at that moment are within $\Delta_{2}(t)$ have been in the $x>0$ region. In other words $h_{2}$ is a strict subset of the histories that crossed $x=0$. We can therefore infer from this consistent set, that the particle passed $x=0$ with certainty (probability unity). Interestingly, while this consistent set gives an indirect way to answer the time-of-arrival question, it actually recovers the answer that one would intuitively expect.